# Foundation Models for Structural Health Monitoring

### v1

@@ Abstract
We showcase the effectiveness of our foundation models using data from three operational viaducts. For AD, we achieve a near-perfect 99.9% accuracy with a monitoring time span of just 15 windows. In contrast, a state-of-the-art method based on Principal Component Analysis (PCA) obtains its first good result (95.03% accuracy) only considering 120 windows. On two different TLE tasks, our models obtain state-of-the-art performance on multiple evaluation metrics (R2 score, MAEand MSE%). On the first benchmark, we achieve an R2 score of 0.97 and 0.85 for light and heavy vehicle traffic, respectively, while the best previous approach stops at 0.91 and 0.84. On the second one, we achieve an R2 score of 0.54 versus the 0.10 of the best existing method.
@@ 1 Introduction
Structural Health Monitoring (SHM) is crucial to ensure the longevity and safety of critical infrastructures such as bridges, highways, and tunnels. In fact, these structures are subject to deterioration due to factors such as ageing, environmental conditions, and traffic loads, posing potential safety hazards and substantial financial burdens for maintenance and repairs.
@@
A typical SHM system comprises a network of sensors responsible for measuring multiple parameters relevant to the current state of the structure as well as its surrounding environment. The most widely used SHM sensing devices include fiber optic sensors [1, 2, 3, 4], accelerometers [5, 6], cameras [7] and strain gauges [8]. Among these options, networks of Micro-Electro-Mechanical Systems (MEMS) for vibration monitoring stand out due to their cost-effectiveness and relatively simple deployment in diverse environments. Data gathered by these accelerometers are then used as the input of a Machine Learning (ML) model to address tasks such as AD or TLE. These devices have demonstrated high accuracy compared to other technologies on several SHM tasks [9].
ML approaches for vibration-based SHM span a diverse range of techniques, from classic methods as in[10, 11], to more recent deep learning solutions such as Convolutional Neural Networks (CNNs) [12] and Autoencoders [13, 14, 15]. While Transformer neural networks have revolutionized the Computer Vision and Natural Language Processing (NLP) fields, in SHM they have only been applied to camera-based monitoring solutions [16], and not yet (to our knowledge) to vibration-based ones. Furthermore, existing literature primarily focuses on models tailored to single tasks, like AD or TLE, and does not consider the possibility of developing a unified, foundation approach capable of addressing multiple SHM challenges with a single ML model.
In this paper, we introduce two key contributions to acceleration-based SHM. We consider for the first time the application of Transformers this task, using a Masked Autoencoder architecture inspired by [17]. Furthermore, building on the recent breakthroughs of foundation models across diverse domains (language, vision, etc..) [18], we explore self-supervised learning to develop models that learn generalizable representations of acceleration data for SHM, and can be then fine-tuned to perform multiple diverse tasks with high performance, including AD and different forms of TLE. This work demonstrates the potential of foundation models for advancing data-driven SHM, especially to cope with the difficulty of collecting large amounts of labelled data, which requires time-consuming and costly instrumentations of the target infrastructure with other types of sensors, such as cameras or Weigh-in-Motion (WiM) [19] systems to obtain ground truths.
@@
  • We consider three different SHM datasets, including a newly collected one, larger than those introduced in previous works and fully labelled.
  • We build a Transformer-based Masked Autoencoder inspired by [17] to test if a foundation model approach could outperform individual task-specific models on SHM tasks. By pre-training on all three datasets without using labels (self-supervised learning) and then fine-tuning on each specific task, we achieve better results than training three separate models from scratch. Notably, although all our datasets are labelled, this 2-stage training protocol only uses supervision signals during fine-tuning, simulating the availability of large amounts of unlabelled samples (much easier to collect in practice) and a relatively smaller amount of labelled ones.
@@ 3.1 Anomaly Detection
Concerning visual AD, in [16], a ViT-like approach is applied to the semantic segmentation of images to identify damaged structural components in a railway viaduct. This approach outperforms existing vision-based methods by a significant margin with 97% and 90% mean Intersection Over Union (mIoU) for component and damage segmentation respectively. [29] proposes a solution for detecting damage on a bridge from displacement measurements using a roving multi-camera vision system. Five unsupervised AD techniques are compared (Autoencoder, K-Nearest Neighbours, Kernel Density, Local Outlier Factor and Isolation Forest), achieving an F1 score that ranges from 0.96 to 0.97 across a series of controlled laboratory tests.
Another line of work [30] proposed a charging scheduling to replenish a wireless rechargeable sensor network for bridge monitoring by an unmanned aerial vehicle (UAV).
However, most AD methods for SHM rely on acceleration data. In [31], the authors use a 1-dimensional CNN to estimate the Probability of Damage, evaluating nine scenarios of increasing damage severity. Similarly, [12] proposes a CNN for multiclass unbalanced AD. They first extract statistical features from acceleration data (i.e. maximum, minimum, mean, variance, skewness, kurtosis), which are then used as the input of 3 stacked convolutional blocks, followed by global average pooling and softmax to predict the category of the damage, choosing from seven distinct anomalies. This method achieves an overall accuracy of $97.6\%$ on a real-world dataset of a long-span cable-stayed bridge in China. In [32, 33], the authors explore spiking neural networks to accurately classify the damage in a viaduct, achieving an accuracy of 95% and Matthew’s correlation coefficient (MCC) values of 0.88.
All aforementioned approaches utilize supervised learning, which requires both healthy and damaged samples for training. In contrast, the authors of [34] employ an Autoencoder for indirect damage diagnosis in a simulated bridge environment, extracting representative features from vehicles’s vibration response, which are then used for an unsupervised damage severity comparison model and a semi-supervised damage severity estimation model. When predicting the damage severity in a range of 30 possible levels, their predicted level is off by $5$ levels on average with respect to the ground truth. In [35], an unsupervised learning method for Recurrent Neural Network (RNN) is employed to detect and identify structural damage, achieving 93% and 85% accuracy for detection and localization. The dataset used for this study was obtained through simulations rather than from a real-world bridge.
Lastly, [36] is the most recent and promising approach for AD on vibration data. It is based on Principal Component Analysis (PCA) compression and does not require labelled anomalies. Specifically, the PCA coefficients are fitted only on healthy data. Then, anomalous vibrations associated with damages, being out-of-distribution samples, will be poorly reconstructed from principal components extracted with those coefficients. Accordingly, this AD scheme classifies all samples with a PCA reconstruction error higher than a threshold as anomalies. On one of the datasets we use, collected from a real-world viaduct located in northern Italy on a state highway, they can reach 98.8% accuracy, 100% specificity, and 97.33% sensitivity on the anomaly detection task. In Section 6, we consider this last method as the state-of-art to compare with.
@@ 3.2 Traffic Load Estimation
Kamkar et al. [7] use smart cameras to detect, classify, and quantify vehicles on public bridges, achieving a notable accuracy of 92.1% in high-traffic scenarios. However, the system’s performance decreased substantially to 74.8% when faced with reduced lighting conditions on the bridge. This finding emphasizes how visual-based methods are sensitive to environmental factors, particularly lighting, which can impact their overall dependability. An alternative method, as illustrated in [37, 38], employs magnetic sensors paired with adaptive thresholds and classification trees. This approach demonstrates varying levels of accuracy, ranging from 84.7% to 99.9%, depending on the traffic conditions analyzed. Notably, in [37], the authors achieve the highest precision in a testing scenario that involves a single vehicle, consistently moving at a slow speed, which does not reflect real-world traffic conditions. In [2, 3, 4], the authors investigate alternative sensor technologies, including fiber optic sensors and infrared sensors, for vehicle detection and classification. These technologies demonstrate the potential to achieve accurate results; however, they may encounter challenges related to sensitivity to environmental factors as well as higher installation costs that could hinder their widespread implementation outside urban environments. For example, deploying single roadside-installed sensor fiber strips might not be feasible in less populated areas. The work of [39] develops an agent that uses TLE information to efficiently control traffic signals based, adjusting speed profiles for approaching vehicles to smooth traffic flows.
Very few works use acceleration data for TLE. In [40], authors showcase the superiority of MEMS accelerometers, proving them to be insensitive to sensor placement and robust to the potential presence of damage in the bridge. In [41], the TLE problem based on SHM sensors was framed as an anomaly detection problem, where vibration patterns that deviated from a periodically updated baseline were associated with the passage of a vehicle. This method is interesting because it does not rely on labelled data, but the overall performance is not sufficient in high-traffic conditions. They use k-means clustering to distinguish 2 groups of vehicles (light and heavy), obtaining in low traffic load conditions extracted features that are significantly different between the two groups, thus resulting in a p-value $<$ 0.0001. [10] is the most recent work on TLE with accelerometer data. It compares several classic machine learning approaches, such as Linear Regressor, Random Forest, k-nearest Neighbors, Support Vector Regressor, and a deep learning approach based on Multi-Layer Perceptron, using data from MEMS accelerometers, achieving state-of-the-art performance with an R2 of $0.91$ on light vehicles and $0.84$ on heavy vehicles. This is the method that we compare with in Section 6.
4 Structural Health Monitoring Installations and Datasets

### v2

@@ Abstract
We showcase the effectiveness of our foundation models using data from three operational viaducts. For AD, we achieve a near-perfect 99.9% accuracy with a monitoring time span of just 15 windows. In contrast, a state-of-the-art method based on Principal Component Analysis (PCA) obtains its first good result (95.03% accuracy), only considering 120 windows. On two different TLE tasks, our models obtain state-of-the-art performance on multiple evaluation metrics (R2 score, MAE% and MSE%). On the first benchmark, we achieve an R2 score of 0.97 and 0.90 for light and heavy vehicle traffic, respectively, while the best previous approach (a Random Forest) stops at 0.91 and 0.84. On the second one, we achieve an R2 score of 0.54 versus the 0.51 of the best competitor method, a Long-Short Term Memory network.
@@ 1 Introduction
Structural Health Monitoring (SHM) is crucial to ensure longevity and safety of critical infrastructures such as bridges, highways, and tunnels. In fact, these structures are subject to deterioration due to factors such as ageing, environmental conditions, and traffic loads, posing potential safety hazards and substantial financial burdens for maintenance and repairs.
@@
A typical SHM system comprises a network of sensors responsible for measuring multiple parameters relevant to the current state of the structure as well as its surrounding environment. The most widely used SHM sensing devices include fiber optic sensors [1, 2], accelerometers [3, 4], cameras [5] and strain gauges [6]. Among these options, networks of Micro-Electro-Mechanical Systems (MEMS) for vibration monitoring stand out due to their cost-effectiveness and relatively simple deployment in diverse environments. Data gathered by these accelerometers are then used as the input of a Machine Learning (ML) model to address tasks such as Anomaly Detection (AD) or TLE. These devices have demonstrated high accuracy compared to other technologies on several SHM tasks [7].
ML approaches for vibration-based SHM span a diverse range of techniques, from classic methods as in[8, 9], to more recent deep learning solutions such as Convolutional Neural Networks (CNNs) [10] and Autoencoders [11, 12]. While Transformer neural networks have revolutionized the Computer Vision and Natural Language Processing (NLP) fields, in SHM they have only been applied to camera-based monitoring solutions [13], and not yet (to our knowledge) to vibration-based ones.
In this paper, we introduce two key contributions in the direction of a new family of models for acceleration-based SHM. First, we explore, for the first time, the application of Transformers to this task. Second, building on the recent breakthroughs of foundation models across diverse domains (language, vision, etc.) [14], we explore self-supervised learning to take advantage of large amounts of easily available unlabelled data and develop general backbone feature extractors, which can be then fine-tuned for multiple SHM tasks. By doing so, we outperform current, single-task ML approaches, which are instead trained on small datasets in a fully supervised way.
More specifically, we adopt a transformer-based Masked Autoencoder architecture inspired by [15], which has become a standard in self-supervised learning, demonstrating better performance compared to classical autoencoder architectures. This approach demonstrates the potential of foundation models for advancing data-driven SHM, particularly in overcoming the challenge of collecting large amounts of labeled data, a process that is typically time-consuming and costly, requiring extensive instrumentation of target infrastructures with additional sensors, such as cameras or Weigh-in-Motion (WiM) systems [16] to obtain ground truth data.
@@
  • We consider three different SHM datasets, including a newly collected one, larger than those introduced in previous works and fully labelled. These datasets are collected by a specialized company with patented sensor networks, expert in real-time SHM. They include two viaducts and a bridge, representative of common Italian infrastructures.
  • We build a Transformer-based Masked Autoencoder, inspired by [15], as the first Foundation Model for SHM, capable of addressing multiple tasks within a unified framework. By pre-training on three diverse datasets without labels (self-supervised learning) and fine-tuning on each task, we outperform individually trained models, demonstrating the benefits of transferable representations. This two-stage training approach simulates the realistic SHM scenario where large unlabelled datasets are easily available, but labelled samples are scarce.
@@
  • We demonstrate the feasibility of deploying our models on a gateway-class edge device, i.e., the NVIDIA Jetson Nano, showing a latency well below the real-time constraint of the three analyzed use cases.
@@ 3.1 Anomaly Detection
Concerning visual AD, [13] applies a ViT-like approach for semantic segmentation to identify damaged structural components in a railway viaduct, achieving 97% and 90% mIoU for component and damage segmentation, respectively, and outperforming existing vision-based methods. Similarly, [26] detects bridge damage using a roving multi-camera system, comparing five unsupervised techniques (Autoencoder, K-Nearest Neighbours, Kernel Density, Local Outlier Factor, and Isolation Forest) and achieving F1 scores between 0.96 and 0.97 in controlled tests. While visual approaches reach impressive accuracy, they incur high installation costs and are effective only under good visibility conditions (e.g., not during night or foggy days).
Therefore, most AD methods for SHM rely on acceleration data. For example, [27] employs a 1-dimensional CNN to estimate the Probability of Damage across nine scenarios of increasing severity. In [10], statistical features (maximum, minimum, mean, variance, skewness, kurtosis) extracted from acceleration data feed into a CNN with three convolutional layers, global average pooling, and softmax, achieving 97.6% accuracy on a long-span cable-stayed bridge dataset. Spiking neural networks have also been explored in [28, 29], reaching 95% accuracy and Matthews Correlation Coefficient (MCC) values of 0.88 for viaduct damage classification. These supervised approaches require extensive labeled data, limiting their scalability. Furthermore, they are specifically trained on a single infrastructure, while their generalization over multiple ones is not tested.
In contrast, unsupervised methods such as [30] use Autoencoders to extract features from vibration responses for indirect damage diagnosis, though predictions can deviate by an average of five levels on a 30-level scale. Similarly, [31] employs an unsupervised Recurrent Neural Network (RNN) for damage detection and localization, achieving 93% and 85% accuracy, respectively, albeit on simulated data. The most promising unsupervised AD method for vibration data is presented in [32], which leverages PCA compression without labeled anomalies. PCA coefficients are fitted solely on healthy data, so anomalous vibrations are poorly reconstructed and flagged when the reconstruction error exceeds a threshold. On one of our real-world datasets—collected from a viaduct in northern Italy—this approach achieved 98.8% accuracy, 100% specificity, and 97.33% sensitivity, and is used as the state-of-the-art benchmark in our comparisons (see Section 6).
@@ 3.2 Traffic Load Estimation
Kamkar et al. [5] employ smart cameras to detect, classify, and quantify vehicles on public bridges, achieving 92.1% accuracy in high-traffic scenarios. However, their performance falls to 74.8% under low-light conditions, highlighting the sensitivity of visual-based methods to environmental factors. In contrast, methods using magnetic sensors with adaptive thresholds and classification trees [33, 34] report accuracy ranging from 84.7% to 99.9%, but they often achieve their best performance under controlled conditions (e.g., single, slow-moving vehicles) that do not reflect real-world traffic. Alternative sensor technologies, such as fiber optic [1] and infrared sensors [2], can also attain high accuracy; however, they also face challenges related to environmental sensitivity and high installation costs (they have to be embedded in the asphalt, requiring road interventions), limiting their feasibility in less populated areas. [35] develops an agent for adjusting traffic signals based on TLE data, but this approach does not directly address vehicle classification or quantification challenges. On the other hand, acceleration-based approaches for TLE are less common. In [36], the superiority of MEMS accelerometers is demonstrated, showing robustness to sensor placement and structural damage. In [37], TLE is framed as an anomaly detection problem, using k-means clustering to differentiate between light and heavy vehicles based on deviations from a periodically updated baseline. Although this method avoids reliance on labeled data, its performance in high-traffic conditions is insufficient. The most recent work, [8], compares several classical machine learning techniques and a deep learning approach (Multi-Layer Perceptron) on MEMS accelerometer data, achieving state-of-the-art R2 scores of 0.91 for light vehicles and 0.84 for heavy vehicles. We use this as the benchmark for comparison in Section 6.
4 Structural Health Monitoring Installations
This section describes the three SHM Use Cases (UCs) considered in our work. The first UC targets anomaly detection, identifying abrupt structural changes, while the other two focus on traffic load estimation with continuous targets. Whereas these three scenarios only regard vehicle communication infrastructures, they serve as a preliminary validation of self-supervised transformer-based models for large-scale SHM systems, which could be extended in the future to other civil structures whose structural health is measured by similar acceleration signals.
4.1 Use Case 1 (UC1): Anomaly Detection Task
This UC utilizes vibration data from a dual-carriageway state highway viaduct in northern Italy. In 2019, a structural strengthening intervention on one of its spans altered the vibration signals. Before the intervention, five SHM nodes were installed under the bridge to monitor its vibrations, providing data from both pre- and post-intervention periods.
We treat post-intervention vibrations as normal and pre-intervention vibrations as anomalous, simulating abrupt structural changes (e.g., due to an earthquake). The sensor nodes use LIS344ALH 3-axis accelerometers [38] (±2/±6g range) and an HTS221 temperature/humidity sensor [39]. Data collection and processing are handled by an STM32L476VGTx microcontroller (ARM Cortex-M4, 80 MHz, 96 KB SRAM, 1 MB Flash).
4.2 Traffic Load Estimation Task
4.2.1 Use Case 2 (UC2)
This UC focuses on a viaduct in Italy, monitored by an SHM sensor network since 2017 (Fig. 2). The viaduct spans  580 m, supported by five concrete pillars, with individual spans ranging from 67 m to 112 m. Each span consists of a steel caisson housing 12 prestressing cables, with cross-section heights varying from 6.0 m at the bearings to 3.0 m at mid-span.
Given that the structural integrity of prestressed bridges depends on cable durability, a sensor system (Fig. 2) was installed to detect breakages. Since September 2017, a network of 90 sensors has been actively collecting and storing data in the cloud. Each sensor node features an STM32F405RG microcontroller (MCU) for data acquisition and optional analytics, alongside the same accelerometer, temperature, and humidity sensors used in Section 4.1.
4.2.2 Use Case 3 (UC3)
This UC examines a reinforced concrete girder bridge with 18 spans, totaling 583 meters in length, with two lanes. The bridge follows an isostatic static scheme, with 20-meter spans, except for the first (10 m) and last (29.5 m) spans. Since 2022, an industrial SHM system has continuously recorded data.
The system includes 282 MEMS biaxial inclinometers and 142 MEMS triaxial accelerometers (±2 g full scale, 100 Hz sample rate), evenly distributed across spans. Accelerometers monitor the two external beams per span, with sensors positioned at the quarter, third, and midspan locations.
Additionally, a Weigh-in-Motion (WiM) system is installed 600 meters before the bridge, providing ground truth traffic data—including lane, vehicle length, weight, speed, and axle count. To align WiM data with SHM sensor readings, a time offset correction is applied based on vehicle speed and distance between the WiM system and each sensor.
5 Methodology
This section presents our novel approach to vibration-based SHM leveraging self-supervised learning and Transformer foundation models. Firstly, we outline the data processing toolchain and train/test splitting (Section 5.1). Then, we delve into our model architecture (Section 5.2) and training strategies and model exploration (Section 5.3, Section 5.4). The core objective of our work is to understand whether it is possible to pre-train a large Transformer model on substantial amounts of unlabeled data and then fine-tune it to realize different SHM tasks. We consider both AD and TLE as downstream tasks to validate this hypothesis, focusing on all three Use Cases described in Sec. 4. Specifically:
The complete pipeline is shown in Figure 3.
5.1 Pre-processing Pipeline
This section outlines the pre-processing applied before inputting data into our masked autoencoder. The pre-processing settings are consistent for training and validation data across all analyzed UCs, except when expressly indicated.
The first pre-processing step is windowing, i.e., the division of the time-series data into time windows, which the algorithm can process. For UC1, we use 5s windows, as in the reference paper [32], with a stride of 2s to increase the number of training samples. For the other use cases, we enlarge the window to 60s since the reference study [8] reported the best performance with this dimension, using a 2s stride in UC2, and 15s stride in UC3, given the extremely lower traffic on the viaduct (with a 2s stride, the information contained in successive windows was almost identical, while causing a linear increase in training time). All windows are composed of uni-dimensional signals since we always employ either a single sensor or multiple sensors, but consider them separately. We also only consider the z-axis of the acceleration data. Thus, each window is always a single vector of dimension $T=100\times T_{win}$, where 100 is the acceleration sampling frequency in all three datasets. All time windows are normalized with mean and standard deviation.
After creating the windows, we apply energy filtering, i.e., we disregard as irrelevant those windows whose cumulative energy is under a certain threshold, $th=3.125\times(10^{-5})$ for UC1 and $th=1.25\times 10^{-6}$ for UC2 and UC3. Energy thresholds are computed using the iterative process proposed in [32].
Lastly, building on the findings of [15], which demonstrated that masked autoencoders achieve superior signal reconstruction in the frequency domain, we adopt a similar strategy in our approach. Specifically, we transform the raw acceleration data into its spectrogram representation, enabling the model to learn meaningful structural patterns through the reconstruction of masked spectrogram patches. Precisely, we feed the model with 100x100 time-frequency representations in the form of spectrograms. The first dimension corresponds to the temporal domain, while the second captures the frequency domain. While we notice that this window is big enough to obtain good task performance, we did not explore the downscaling of the spectrogram image to reduce the models’ computational complexity, as the reduction of the input size has an impact only on the dimension of the last layer, whose size is negligible with respect to the whole model size.
$y=\frac{\sum_{t=1}^{T}(l_{t}=k)}{10}$ (1)
5.2 Foundation Model Architecture
Figure 4: Architecture of our foundation model transformer-based masked autoencoder during pre-training.
Our model architecture draws inspiration from the Masked Autoencoder proposed in [15], which was originally introduced to learn self-supervised representations from audio spectrograms. The model uses a stack of standard Transformer blocks for both the encoder and the decoder, each consisting of multiple attention layers. As a backbone, we use ViT-Base (ViT-B). The model takes as input acceleration spectrograms divided into non-overlapped regular grid patches, with a configurable portion of them being masked during training. The complete architecture is shown in Fig. 4.
Architectural modifications are necessary to fine-tune our foundation model based on the downstream tasks considered. Namely, the model architecture remains unchanged when fine-tuning on UC1, because the task is framed as an outlier detection [32]: during training, the model learns the latent representation of normal data only, so that during testing, the decoder is expected to fail in accurately reconstructing data that deviates from its learned patterns. Accordingly, those samples whose reconstruction error (computed as detailed in Sec. 5.3.1) exceeds a user-defined threshold are classified as anomalies or outliers. On the other hand, the decoder is removed when fine-tuning on TLE tasks (UC2 and UC3), and a fully connected layer with a single neuron is appended directly to the latent representation produced by the encoder to produce the desired continuous traffic output defined in Eq. 1. The encoder is initialized with the weights learned during the self-supervised pre-training step, ensuring the transfer of the learned, general feature extraction. Additionally, to handle potential differences in input resolution, we interpolate the position embeddings as done in DeiT [40], preserving spatial consistency in learned representations.
5.3 Training Phases
5.3.1 Pre-training
The full encoder-decoder model, with the architecture of Fig. 4 is initially pre-trained on a large amount of unlabeled data. We select pre-training data following the split described in Section 2, combining the training sets of all three UCs. Note that, although our data have labels, we do not use them in this phase, to simulate a scenario where the costly process of labelling all available samples is not affordable.
During pre-training, the non-overlapping spectrogram patches of each input sample are masked with probability $p$ (masking ratio) and aligned, before being passed to the transformer encoder. The latter then processes the non-masked patches to produce latent representations. Encoded patches are padded with trainable mask tokens and fed to the decoder, whose objective is to reconstruct the original spectrogram, including the missing parts. The encoder and decoder are jointly optimized to minimize the Mean Squared Error (MSE) between the reconstruction and the input spectrogram, averaged over masked patches. This pre-training is completely self-supervised and task-independent. Its objective is to learn robust latent representations of SHM data that allow the model to infer the underlying patterns in the data despite the missing information.
5.3.2 Fine-tuning
The details of the fine-tuning phase differ depending on the downstream task. For UC1 (AD), as mentioned above, we maintain the autoencoder structure of our model. Accordingly, the model is fine-tuned in an unsupervised way, using the same procedure just described, but considering only normal data from the UC1 dataset. At test time, we use a threshold on the reconstruction error to identify anomalies. For both our models and the baseline PCA, we follow a procedure similar to [32] to determine it. Namely, the threshold is initialized to the average MSE on non-anomalous training data, and then progressively increased by a small step (1/1000th of the average), until 100% specificity is achieved on a validation set, obtained as a 20% sample of the training dataset.
For UC2, the pre-trained model is fine-tuned on acceleration data and video recording on the viaduct presented in Sec. 6.1. Two separate fine-tunings are executed for light and heavy vehicle TLEs, starting from the same pre-trained model. For UC3, we consider the first half of the second day’s data to fine-tune our foundation model, as described in 6.1. For these two downstream tasks, we follow a classic supervised fine-tuning procedure, where the model learns to predict the traffic load values rather than to reconstruct the input spectrograms. To fine-tune the encoder weights and train the added fully-connected layer, we use an initial lower learning rate of $0.25\cdot 10^{-5}$ that is updated with a half-cycle cosine decay schedule. As a loss function, we use the MSE between the newly inserted fully connected layer’s output prediction and the ground truth target variable.
5.4 Model Exploration and Knowledge Distillation
To select our model’s hyperparameters, we do a search over the encoder/decoder depth and heads, the embedding dimensions, and the masking probability $p$ of the masked autoencoder. To make this search manageable, we sample 20% of the pre-training data, further splitting them into 80% for training and 20% for validation. The task is the same as the pre-training, i.e., reconstructing the masked input data. The best configuration is selected based on the lowest error achieved on the validation set. We employ Bayesian optimization with the open-source Optuna framework [^1] https://optuna.org, fixing $p=0$. Then, we explore $p$ through a grid search with a step of 0.2. Table 1 summarizes the search space, with the optimal configuration highlighted in bold.
Table 1: Explored space of architectural hyperparameters for the Transformer encoder-decoder
Parameters Values explored
Encoder depth $1,3,5,\textbf{7}]$
# Encoder Heads $\textbf{8},12,16]$
Encoder embedding $1536,\textbf{768},384,192]$
Decoder depth $\textbf{2},4,6,8]$
# Decoder Heads $8,\textbf{16},24]$
Decoder embedding $1024,\textbf{512},256,128]$
Masking ratio $0,0.2,0.4,0.6,\textbf{0.8}]$
After this step, we also define a family of smaller models to explore the possibility of reducing the computational burden with minimal accuracy loss, targeting the deployment on sensor nodes with an extremely low memory budget. Starting from the frozen configuration of Table 1, we systematically reduce the embedding dimensions of both the encoder and decoder, starting from the baseline configuration $(e_{dim},d_{dim})=(768,512)$ and scaling them down by a factor of $2$ until reaching $(e_{dim},d_{dim})=(24,16)$. These new models are pre-trained on all datasets and subsequently fine-tuned on the three downstream tasks, following the same procedure described in the previous section. Additionally, to inherit our big model performance, we considered a modified fine-tuning procedure enhanced with Knowledge Distillation (KD). In this setup, considering the baseline model as a teacher, the student utilizes a multi-objective fine-tuning loss function composed of two contributions: the first one, $\mathcal{L}_{task}$, is the MAE computed on student prediction against the ground truth, while the second one, $\mathcal{L}_{KD}$, is the Root Mean Squared Error (RMSE) between student and teacher predictions. Through an empirical approach, we assessed that evenly balancing these two contributions leads to faster convergence of the overall loss, whose function becomes:
6 Results